OpenStudy (anonymous):

What is the simplest form of the expression 14ab^3 /7a^-2b^-1

OpenStudy (jdoe0001):

\(\bf \cfrac{14ab^3}{7a^{-2}b^{-1}}\implies \cfrac{14ab^3}{7}\cdot \cfrac{1}{a^{-2}}\cdot \cfrac{1}{b^{-1}} \\ \quad \\ \textit{keep in mind that }\large \cfrac{1}{a^{-\color{red} n}}\implies \cfrac{1}{\frac{1}{a^{\color{red} n}}}\implies a^{{\color{red} n}}\)

OpenStudy (anonymous):

I'm not understanding a single thing you just said

OpenStudy (anonymous):

\[ \frac{14ab^3}{7a^{-2}b^{-1}} \times \frac{a^2}{a^2}= \frac{14a^{1+2}b^3}{7a^{-2+2}b^{-1}} = \frac{14a^{3}b^3}{7b^{-1}} \]

OpenStudy (anonymous):

Next step is multiply the top and bottom by \(b\).